### v2

# Software Framework for Optically Accessible Quantum Memory Using Group-IV Color Centers in Diamond

Abstract
In the rapidly evolving field of quantum technology, the precise and detailed description of quantum components is not just a necessity but the foundation for advancing research, development, and applications. Optically accessible quantum memories are key building blocks for devices such as quantum repeaters and two-factor authentication. The memory we describe here is based on a tin-vacancy color center coupled to a highly efficient cavity. It leverages state-dependent reflection from the cavity and implements high-fidelity fractional single qubit gates via a train of optical $\pi/8$ pulses. We also describe its operation under microwave control, further extending our analysis. Our primary contribution in this work is the integration of this device model into a standardized software framework for quantum memory architectures.
Key words: quantum communication, quantum optics, quantum control
I Introduction
Figure 1: a) A time-bin qubit $|\psi\rangle$ is stored in a Sawfish [1, 2] spin-photon interface that hosts a G4V. Storage is achieved via the electronic spin. Fast switches (S) are deployed throughout the circuit to enable efficient measurement and routing of the photonic state. The $X$-basis measurement is performed using a fast switch integrated with an imbalanced Mach-Zehnder interferometer. The storage procedure is completed once the photonic qubit is measured in the $X$-basis. For retrieval, the spin state is read out by entangling it with a photon from an extra single-photon source (SPS) after a time-bin preparation state (TBPS) [3, 4, 5]. A subsequent $Z$-basis measurement on the spin is performed. b) Using the SnV center as a G4V, spin-photon entanglement is mediated through a Sawfish cavity-to-fiber interface [1, 2]. The SnV is modeled as a four-level system characterized by its spontaneous emission rates $\gamma_{1A},\gamma_{2B}$, spectral contrast $\Delta\omega_{s}=\omega_{2B}-\omega_{1A}$ spin splitting $\omega_{s}$, and atomic transition frequencies $\omega_{1A},\omega_{2B}$, while interacting with a cavity mode $a_{c}$ at frequency $\omega_{c}$ for the spin-dependent reflection steps. c) Entanglement is generated employing a reflection scheme containing spin-dependent reflection steps and a $\pi/2$ rotation where the spin is initialized in the state $|1\rangle$. The $\pi/2$ rotation is applied to the spin before the final reflection event. This rotation can be implemented either using a sequence of four $\pi/8$ optical Raman pulses or utilizing microwave control. Figures b) and c) are adopted from [6].
Optically accessible quantum memories become of increasing interest in the era of quantum communication [7, 8, 9], computing [10, 11, 12] and sensing [13, 14, 15]. Below, we highlight key applications: First generation quantum repeaters, a key component in scalable quantum networks, rely on a quantum memory to extend entanglement over long distances [16, 17, 18]. Beyond enabling long-distance entanglement, quantum memories also play a crucial role in quantum security applications [19] such as quantum authentication [20, 6]. While quantum memories are important for quantum security and communication they are also highly relevant for scalable quantum computing [21], and especially for the development of a quantum random-access-memory (QRAM) [22]. Finally, the field of quantum sensing benefits from the use of quantum memories since they can help in increasing the sensor’s sensitivity [23].
There are various types of quantum memories: The atomic-ensemble based quantum memories can map a photonic qubit onto a collective-spin excitation shared between two atomic ensembles [24], or onto a spatially multiplexed cold atomic ensemble [25] achieving storage times in the range of tens of microseconds. Another type of memory is based on solid-state systems: Prominent examples are rare-earth ion-doped crystals [26] or color centers in diamond [27, 28], which are able to store quantum information using methods such as the atomic frequency comb (AFC) protocol [26, 29] or by using a spin-photon interface and long lived nuclear spins in the diamond lattice [27]. Superconducting quantum memories are also being actively researched [30, 31, 32, 33], which can be readily integrated into a superconducting quantum computing platform. Superconducting qubits are so far not optically accessible and require superconducting waveguides for networking [34]. Finally, hybrid quantum memories combine physically distinct systems to push the scheme’s performance beyond that of the individual components. Take for example storage and retrieval of photons from a quantum dot single-photon source interfaced with an atomic quantum memory [35], based on the Off-Resonant Cascaded Absorption (ORCA) protocol [36].
In this work, we focus on a hybrid quantum memory scheme that uses single-photon sources [37], such as quantum dots [35], solid state emitter [38] or heralded single photons from spontaneous parametric down conversion [39], together with an atom-like memory system based on negatively charged group-IV color centers in diamond (G4V) [40]. The latter is integrated into a single-sided nanophotonic cavity, specifically the Sawfish cavity [1, 2]. Tin-vacancy centers (SnVs) are a particularly suitable choice for such quantum memories [28, 41, 42, 43] when integrated into nanophotonic structures. Due to their reduced sensitivity to charge noise [44], their optical coherence properties are largely preserved. By using state‑dependent reflection from the cavity and implementing high-fidelity fractional quantum gates through a pulse train of optical $\pi/8$ pulses, our specific memory design can achieve gate fidelities exceeding 99% depending on the bandwidth. We also consider microwave control for the memory scheme, which extends the viability to longer storage times, at the cost of reduced operational rates [45].
The goal of this work is to present the technical background for describing the device within a standardized software framework for quantum memory architectures [46]. The software module we developed allows users to compute Kraus operators for the read-in and read-out processes under optical or microwave control, respectively, and to specify photon-generation fidelity, bandwidth, and the nanophotonic system’s temperature. When microwave control is selected, users may also adjust strain as well as DC and AC magnetic field strengths and orientations. To support this module, we provide a detailed description of the model for storing and retrieving a time-bin photonic qubit in a spin state, which serves as the basis of the software component.
To model the spin–photon interface, we use closed-form integral expressions covering detailed modeling of spin-photon interaction for the read-in process, accounting for both an imperfect and broadband photon source and imperfect spin rotations, which are generated by optical control [6, 47] or by microwave control [45], depending on the color center’s temperature and strain. For the read-out process, we assume an ideal single-photon source but include imperfect spin rotations.
The standardized description of the memory requires the use of Kraus-operators [48], which we derive for both: the read-in and read-out process. Such a Kraus representation of the quantum memory system offers the advantage of seamless integration for benchmarking quantum repeater performance and various quantum information applications making it an ideal technical contribution for a large quantum memory simulation software. In [46] we provide the corresponding module.
Sec. II introduces our proposed hybrid quantum memory scheme consisting of a single-photon source and a group-IV color center in diamond. Sec. III outlines the mathematical principles concerning the described system. Sec. IV provides numerical examples for the Kraus-operators. Finally, Sec. V summarizes the results and provides an outlook.
II Quantum Memory Scheme
The quantum memory scheme is illustrated in Figure 1. A single-photon source emits photons which are transmitted via a fiber to the quantum memory formed by a highly efficient single-sided Sawfish crystal cavity [1, 6] (see Figure 1b). These photons are entangled with the cavity’s spin and subsequently the photon is measured in the $X$-basis, completing the read-in process. For the read-out an extra single-photon source (SPS) is required. The respective photon from the SPS is entangled with the saved spin of the color center. Afterwards the spin is measured in the $Z$-basis. The reflection scheme for the read-in and read-out process contains an early reflection, a $\pi/2$ rotation, a late reflection and a measurement as illustrated in Figure 1a,c. The reflection steps realize a controlled phase gate which can be optimized by tuning the cavity loss rate $\kappa$, mode frequency $\omega_{c}$ and central frequency of the emitter $\omega_{0}$ depending on the bandwidth $\gamma$ of the photon. The spin rotation relies on coherent control of the electron spin of the SnV. That is done using optical [6] or microwave control [45]. Optical control is performed using a Raman scheme [6] applying a train of optical $\pi/8$ pulses to achieve a high fidelity $\pi/2$ rotation encountering photonic and phononic processes. The subsequent $X$-basis measurement and thereby the overall saved state after the read-in process represents an important set of the relevant error channels: depolarization during photon generation, broadband photons and control imperfections due to photonic and phononic processes inside the SnV [6]. Beyond [49] we take crosstalk during the spin-photon interaction into account.
The read-out process uses an extra single-photon source whose photon gets entangled with the spin using the reflection scheme with a subsequent $Z$-measurement of the spin. In this work, we make the following key assumptions, which are fixed parameters in the memory software implementation.
Single-photon source: We assume, that the SPSs photons have a fixed bandwidth and model the spin system as a four-level system. We assume a photon generation fidelity $F$, an incoming photonic qubit $|\psi\rangle=\alpha|e\rangle+\beta|l\rangle$ for some $(\alpha,\beta)$ with $|\alpha|^{2}+|\beta|^{2}=1$ and bandwidth $\gamma$. For the extra single-photon source we assume narrowband photons with $\alpha=\beta=1/\sqrt{2}$ with perfect fidelity, ensuring no limitations arising from the auxiliary photon source.
Magnetic field: For the presented example calculation, we assume a static magnetic field strength $B_{\rm dc}=3.0$ T and magnetic field orientation $\theta_{\rm dc}=43.11\deg$ for performing optical control. In future iterations of our package we will include variable field strength and angle. For microwave control the static magnetic field strength $B_{\rm dc}$ as well as the magnitude of the microwave control field $B_{\rm ac}$ and their polar angles $\theta_{\rm dc}$ and $\theta_{\rm ac}$ are variable. We assume the azimuthal angles $\phi_{\rm dc}=0$ and $\phi_{\rm ac}=-\pi/2$ for achieving a rotation around the $y$-axis in the Bloch-sphere [45] (the orientations are visualized in Figure 1b.)
Strain: The SnV is in a low-strain environment for optical control and a strained one for microwave control with the compression $E_{x}$ and shear strain $\epsilon_{xy}$. These strain components can be freely chosen for microwave control.
Temperature: The nanophotonic system including the SnV is assumed to be cryogenically cooled to temperatures ranging from $100$ mK to $4$ K.
Power: The average power consumption of the two lasers is calculated over the full-width at half maximum (FWHM) here,
$P_{i}^{L}=W_{i}/(\sigma\sqrt{e}),\quad W_{i}=1/2\epsilon_{0}E_{i}^{2}\lambda_{i}^{3}/(2n^{3})$ (1)
where $\sigma$ is the standard deviation of the laser pulse, $\epsilon_{0}$ is the dielectricity constant, $E_{i}$ for $i=1,2$ describes the laser field’s amplitude, $\lambda_{i}$ is the wavelength of the respective laser and $n$ is the refractive index of diamond. The microwave power is
$P^{\rm MW}=B_{\rm ac}^{2}/2\mu_{0}A^{\rm MW},\quad A^{\rm MW}=(\lambda^{\rm MW}/2n)^{2}$ (2)
where $B_{\rm ac}$ is the magnetic field strength of the driving field, $A^{\rm MW}$ the effective area and $\mu_{0}$ the magnetic susceptibility.
Time: The total processing time is given by
$T_{1}$ $=T_{\rm read-in}+T_{\rm read-out}+T_{s},$ (3)
$T_{k}$ $=T_{\rm tb,k}+T_{g}+T_{m}+T_{c,\rm k}$ (4)
for $k=\rm read-in,read-out$ where $T_{\rm tb,k}$ is the time allocated for a time-bin qubit, $T_{g}$ is the control gate duration, $T_{m}$ is the measurement time, $T_{s}$ is the storage time, and $T_{c,k}$ is the transmission time for the respective process. We set $T_{\rm tb,k}=20T_{\rm lt,k}$, where $T_{\rm lt,k}$ is the lifetime of the single-photon source used for the read-in or read-out, respectively, and choose $T_{g}=40\sigma$ for optical control with $\sigma=\tau_{\pi/8}/2\sqrt{2\ln(2)}$, where $\tau_{\pi/8}=88.33$ ps is the full-width at half maximum of the optimized $\pi/8$ pulse for $B_{\rm dc}=3.0$ T. We take $T_{m}=100$ ps [50], corresponding to the dead time of the photon detectors [51]. The communication time is given by $T_{c,k}=L_{k}/c$, where $L_{k}$ is the fiber length for read-in or read-out, respectively and $c$ is the speed of light in the fiber.
III Modeling
We outline the steps to compute Kraus-operators that map the photonic state to the spin state during the read-in process. We assume the spin is initialized in the state $|1\rangle$. We then compute Kraus-operators mapping the stored spin state back to the photonic qubit for the read-out process.
III.1 Single-Photon Source
Photons from single-photon sources can be modeled in time-bin encoding with the state space
$\mathcal{H}={\rm span}\{|e\rangle,|l\rangle\}.$ (5)
We define the time-bin qubits as superpositions of single photons $|\omega_{e,l}\rangle$ in frequency space so that
$|e,l\rangle=\int_{\mathbb{R}}S(\omega-\omega_{0})|\omega_{e,l}\rangle\,{\rm d}\omega$ (6)
where $S(\omega)$ is the emission spectrum with a central frequency $\omega_{0}$. We assume a Lorentzian intensity profile [52], i.e.
$S(\omega)=\mathcal{N}\frac{\epsilon_{0}}{{\rm i}\omega+\gamma/2}$ (7)
with the normalization factor $\mathcal{N}=\left(\int_{\mathbb{R}}|S(\omega)|^{2}\right)^{-1/2}$ and the bandwidth $\gamma$.
An arbitrary photonic qubit is then given by
$|\psi_{\rm ph}\rangle=\alpha|e\rangle+\beta|l\rangle,\quad|\alpha|^{2}+|\beta|^{2}=1.$ (8)
We assume that the emitted photon gets depolarized during its generation according to the quantum channel [53]
$\mathcal{E}(\rho_{\rm ph})=(1-\epsilon)\rho_{\rm ph}+\epsilon{\rm tr}(\rho_{\rm ph})\frac{\mathds{1}}{2}$ (9)
with $\epsilon=2(1-F)$ and $\rho_{\rm ph}=|\psi_{\rm ph}\rangle\langle\psi_{\rm ph}|$ where the quantity $F$ denotes the mixed state fidelity between the states $\rho_{\rm ph}$ and $\mathcal{E}(\rho_{\rm ph})$ [54] .
III.2 Cavity Design
To store a photon with a finite bandwidth in the cavity coupled spin we must carefully engineer the cavity such that the controlled phase gate fidelity between the spin and photon [54]
$F_{\rm sp}=\langle\rm Bell|\rho_{\rm sp}|\rm Bell\rangle$ (10)
with $|\rm Bell\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ is maximized (see App. B). To evaluate the fidelity given in Eq. (82), the spin-dependent reflection coefficients are required. We assume that a single cavity mode is coupled to the $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ transitions. The reflection coefficients are given by [55, 49]
$R_{1}(\omega)$ $=-1+\frac{2\kappa({\rm i}\Delta_{1A}+\gamma_{1A}/2)}{({\rm i}\Delta_{c}+\kappa)({\rm i}\Delta_{1A}+\gamma_{1A}/2)+|g_{1A}|^{2}},$ (11)
$R_{2}(\omega)$ $=-1+\frac{2\kappa({\rm i}\Delta_{2B}+\gamma_{2B}/2)}{({\rm i}\Delta_{c}+\kappa)({\rm i}\Delta_{2B}+\gamma_{2B}/2)+|g_{2B}|^{2}}$ (12)
with the cavity loss rate $\kappa$, the detunings $\Delta_{1A}:=\omega-\omega_{1A},\Delta_{2B}=\omega-\omega_{2B}$, the transition frequencies $\omega_{1A},\omega_{2B}$, the atomic decay rates $\gamma_{1A},\gamma_{2B}$, the detuning $\Delta_{c}=\omega-\omega_{c}$ with the cavity mode central frequency of the SnV $\omega_{c}$ and the coupling strength between the SnV and the photonic mode in the cavity
$g_{1A}={\rm i}\sqrt{\frac{\omega_{c}}{2\hbar\epsilon_{0}\epsilon_{r}V}}\langle 1|\bm{\epsilon}\cdot\mathbf{d}|A\rangle,$ (13)
$g_{2B}={\rm i}\sqrt{\frac{\omega_{c}}{2\hbar\epsilon_{0}\epsilon_{r}V}}\langle 2|\bm{\epsilon}\cdot\mathbf{d}|B\rangle~,$ (14)
where the mode volume is $V=V_{\rm eff}\frac{\lambda^{3}}{2n^{3}}$ [56], the wavelength of the cavity mode $\lambda=2\pi c/\omega_{c}$, the speed of light $c$ and the mode orientation $\bm{\epsilon}$. We assume the effective mode volume $V_{\rm eff}=1.8$ and mode orientation $\bm{\epsilon}=\bm{e_{z}}$, i.e. the mode is parallel to the SnV’s symmetry axis [6].
To maximize the fidelity shown in Eq. (82) the software package can optimize the triple $(\omega_{0},\omega_{c},\kappa)$ using simplicial homology global optimization [57, 58]. Optimized values for $(\omega_{c},\kappa)$ yield a cooperativities $\mathbf{C}=(C_{1A},C_{2A},C_{1B},C_{2B})$ calculated by $C_{kl}=2|g_{kl}|^{2}/\kappa\gamma_{kl}$ for $k=1,2$, $l=A,B$, which is necessary for simulating the influence of the Purcell enhancement [59].
III.3 Spin Rotation
The spin $\pi/2$ rotation is an essential step for the quantum memory scheme. It can be performed using optical or microwave control. Depending on the choice of the control the state space is $\mathcal{L^{\rm opt}}$ for optical control and $\mathcal{L^{\rm mw}}$ for microwave control. They are defined by
$\mathcal{L^{\rm opt}}=\{|i\rangle\langle j||i,j=1,...,8\},$ (15)
$\mathcal{L^{\rm mw}}=\{|i\rangle\langle j||i,j=1,...,4\}.$ (16)
Raman control: Our software package propagates $\rho_{ij,0}=|i\rangle\langle j|$ for $i,j=1,2$ subject to the Lindblad-master equation, i.e. [47, 6]
$\dot{\rho}$ $=-{\rm i}[H(t),\rho]+\mathcal{L}_{D}^{(T,\mathbf{C},\bm{B_{\rm dc}})}(\rho),$ (17)
$\rho(0)$ $=\rho_{0,ij}$ (18)
with the Hamiltonian
$H(t)=H_{0}-\bm{\mu}\cdot\mathbf{E}(t),$ (19)
and the dissipator
$\mathcal{L}_{D}^{(T,\mathbf{C},\bm{B_{\rm dc}})}(\rho)=\sum_{k}\gamma_{k}\left(L_{k}\rho(t)L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho(t)\}\right)$ (20)
with the free contribution $H_{0}$ and the light-matter interaction term $\bm{\mu}\cdot\mathbf{E}(t)$ using the optimized parameters for $B_{\rm dc}=3.0$ T and $\theta_{\rm dc}=43.11$ deg as demonstrated in [6]. The detailed description of the Hamiltonian $H^{\rm opt}(t)$ and temperature, cooperativity and magnetic field dependent relaxation rates $\gamma_{k}$ as well as Lindblad-operators $\{L_{k}\}$ modeling photonic decay processes using Fermi’s golden rule and fast phononic processes are explained in [47, 6].
Microwave control: We simulate the $\pi/2$ rotation using microwave control using the Lindblad-master equation shown in Eq. (17) with the Hamiltonian [45]
$H(t)=H_{\rm dc}+H_{\rm ac}(t)$ (21)
consisting of the time-invariant part $H_{\rm dc}$ due to the DC magnetic field and the driving term $H_{\rm ac}(t)$ modeling an oscillating AC magnetic field and the dissipator $\mathcal{L}_{D}^{(T,E_{x},\epsilon_{xy},\bm{B_{\rm dc}})}(\rho)$ containing only the temperature, strain and field $\bm{B_{\rm dc}}$ dependent phononic processes. The $\pi/2$ rotation is achieved by driving the AC field for a quarter of a full Rabi oscillation [45].
III.4 Kraus-Operators
To model the quantum channel for the read-in and read-out we explain the formalism for computing Kraus-operators below. To model a quantum channel $\mathcal{D}$ with Kraus-operators $\{K_{m}\}$ with $\sum_{m=1}^{4}K_{m}^{\dagger}K_{m}\leq\mathds{1}$ [60, 61] such that
$\rho^{ij}(T)=\sum_{m=1}^{4}K_{m}|i\rangle\langle j|K_{m}^{\dagger},\quad i,j=1,2$ (22)
we apply the Choi-Jamiołkowski-Isomorphism [48]. The formalism requires the Choi matrix which is of the form
$J=\sum_{i,j=1}^{2}\mathcal{D}(|i\rangle\langle j|)\otimes|i\rangle\langle j|.$ (23)
The Kraus-operators are
$K_{m}=\sqrt{\lambda_{m}}\begin{pmatrix}\langle 1|\psi_{m}\rangle&\langle 2|\psi_{m}\rangle\\ \langle 3|\psi_{m}\rangle&\langle 4|\psi_{m}\rangle\end{pmatrix}$ (24)
with eigenvalues $\{\lambda_{m}\}$ and eigenvectors $\{|\psi_{m}\rangle\}$ of the Choi matrix $J$. The Kraus map is not trace-preserving because the read-in and read-out process require a measurement.
To evaluate Kraus-operators for the read-in and read-out process we need to propagate the full basis of photonic - and spin states subject to the quantum channel for the read-in and read-out process, respectively. A detailed mathematical exploration of the spin states after measurement including imperfect spin gates using microwave or optical control encountering photonic and fast phononic processes and a detailed spin-photon interaction model including crosstalk which goes beyond the work in [49] is outlined in the Apps. A, C. Let’s assume a photonic qubit $\rho_{\rm ph}$ is passing all reflection steps including the $X$-measurement resulting in the measurement outcomes $\rho_{+}$ and $\rho_{-}$.
To compute Kraus-operators for the read-in process we evaluate
$\mathcal{D}_{\rm read-in}(|i\rangle\langle j|)=\rho_{\rm sp}^{ij}(T_{\rm read-in}),\quad i,j=e,l$ (25)
with the quantum channel for the read-in $\mathcal{D}_{\rm read-in}$ assuming a measurement in $|+\rangle$ for the photonic qubit.
To compute Kraus-operators for the read-out process we evaluate
$\mathcal{D}_{\rm read-out}(|i\rangle\langle j|)=\rho_{\rm ph}^{ij}(T_{\rm read-out}),\quad i,j=1,2$ (26)
with the quantum channel for the read-out $\mathcal{D}_{\rm read-out}$ assuming a measurement in the state $|1\rangle$ for the spin qubit. Details regarding the read-out process are given in the App. D. We compute Kraus-operators by evaluating Eq. (24). For a photonic qubit of the form shown in Eq. (9) the spin state after read-in is
$\rho_{\rm sp}=\frac{\mathcal{D}_{\rm read-in}(\rho_{\rm ph})}{{\rm tr}(\mathcal{D}_{\rm read-in}(\rho_{\rm ph}))}$ (27)
and after read-out we get
$\rho_{\rm ph}=\frac{\mathcal{D}_{\rm read-out}(\rho_{\rm sp})}{{\rm tr}(\mathcal{D}_{\rm read-out}(\rho_{\rm sp}))}.$ (28)
III.5 Approximation Error
Consider the initial state $\rho_{0}=|1\rangle\langle 1|$ and the propagated state subject to the $\pi/8$ rotation at final time $T$ which is called $\rho(T)$. Now define $\tilde{\rho}(T)=\sum_{i,j}\tilde{\rho}_{ij}(T)|i\rangle\langle j|$ with
$\tilde{\rho}_{ij}(T)=\begin{cases}\rho_{ij}(T),\quad{\rm if}\quad i,j=1,2,\\ 0,\quad{\rm else}\end{cases}.$ (29)
Subsequently, the approximation error is
$e=\left\lVert\rho(T)-\tilde{\rho(T)}\right\rVert_{1}$ (30)
where $\left\lVert A\right\rVert_{1}=\max_{1\leq j\leq 8}\left(\sum_{i=1}^{8}|A_{ij}|\right)$ denotes the 1-norm of a matrix. We also use Eq. (29) for evaluating the approximation error for microwave control.
IV Numerical Examples
[…]
V Discussion and Outlook
Our executable quantum memory component leverages tin-vacancy (G4V) centers to store photonic time-bin qubits and demonstrates both robustness and versatility. We optimize $\pi/2$ pulse sequences using either microwave or optical control and tune cavity parameters to simulate the spin–photon entanglement that underlies the read-in and read-out operations. From these simulations we derive the corresponding Kraus-operators, thereby achieving the component’s central objective. The full module is available in [46]. Users may select optical or microwave control for both read-in and read-out, and in all cases they can configure photon-generation fidelity, bandwidth and the nanophotonic system’s temperature. If microwave control is chosen, strain as well as DC and AC magnetic field strengths and orientations are also adjustable; under optical control these parameters remain fixed.
The module can be further extended to enhance its versatility. For example, lifting the current constraints on optical control to permit arbitrary magnetic field strengths would broaden the accessible parameter space. Other group-IV vacancy centers, such as the silicon, germanium, and lead vacancies [62], can also be included, serving an even broader research community.
Overall, the software component [46], which is openly accessible, will enable a broad user base to study group-IV vacancy centers in a quantum-network setting making it a highly valuable research tool.
Author Contributions
Y.S. and G.P. conceptualized the described software module with Y.S. implementing and testing core functionalities. M.B. provided the code specific to microwave spin control. T.S. and G.P developed the core idea and provided overall project supervision. All authors contributed to writing and refining the manuscript.
Conflict of Interest
The authors declare no financial or commercial conflict of interest.
Acknowledgements
We would like to thank Elizabeth Robertson, Fenglei Gu and Johannes Borregaard for fruitful discussions during the development of the model. Funding for this project was provided by the German Federal Ministry of Education and Research (BMBF, project QPIS, No. 16KISQ032K; project DINOQUANT 13N14921, ERC StG project QUREP of the EC, No. 851810).
Author Contributions
Y.S. and G.P. conceptualized the research, with Y.S. conducting the simulations and data analysis. M.B. performed the simulations specific to microwave spin control. G.P. and T.S. developed the core idea and provided overall project supervision. All authors contributed to writing and refining the manuscript.
APPENDIX A SPIN-PHOTON INTERACTION
Group-IV color centers (G4V) in diamond have four states $\{|k\rangle\}$ $k\in G=\{1,2,3,4\}$ in the ground state manifold and four states $\{|l\rangle\}$ $l\in E=\{A,B,C,D\}$ in the excited state manifold. Therefore, the Hamiltonian of the G4V coupled to a single cavity mode is given by
$H=\sum_{m=2}^{8}\epsilon_{m}|m\rangle\langle m|+\omega_{c}a_{c}^{\dagger}a_{c}+\left(\sum_{k\in G,l\in E}g_{kl}\sigma_{kl}+g_{kl}^{*}\sigma_{lk}\right)\left(a+a^{\dagger}\right)$ (31)
where $\epsilon_{m}$ is the bare energy of quantum state $|m\rangle$ with $m\in G\cup E$, $\omega_{c}$ is the central frequency of the cavity mode, and $\sigma_{kl}=|k\rangle\langle l|$ is the transition operator. Here,
$g_{kl}={\rm i}\sqrt{\frac{\omega_{c}}{2\hbar\epsilon_{0}\epsilon_{r}V}}\langle k|\bm{\epsilon}\cdot\mathbf{d}|l\rangle$ (32)
is the coupling strength between $|k\rangle\leftrightarrow|l\rangle$ transition and the cavity mode, where $\mathbf{\epsilon}$ is the cavity mode orientation which is made parallel to the defect’s symmetry axis, i.e. $\bm{\epsilon}=1/\sqrt{3}(1,1,1)$ in the diamond lattice coordinate system, $\mathbf{d}$ is the G4V dipole moment operator, $\epsilon_{r}=5.7$ is the relative permittivity of diamond [63], $V=V_{{\rm eff}}\frac{\lambda^{3}}{2n^{3}}$ is the mode volume, where $V_{{\rm eff}}=1.8$ [64], $n=2.417$ is the refractive index of diamond [56] and $\lambda=2\pi c/\omega_{c}$ is the wavelength of the cavity mode, where $c$ is the speed of the light.
Here, the cavity-G4V setup is considered as an open system with Lindblad operators
$L_{1}=\sqrt{\gamma_{1A}}\sigma_{1A}\;,$ (33)
$L_{2}=\sqrt{\gamma_{1B}}\sigma_{1B}\;,$ (34)
$L_{3}=\sqrt{\gamma_{2A}}\sigma_{2A}\;,$ (35)
$L_{4}=\sqrt{\gamma_{2B}}\sigma_{2B}\;,$ (36)
$L_{5}=\sqrt{2\kappa}a\;,$ (37)
where $\kappa$ is the photon loss rate of the single-sided cavity (a half-width at half-maximum of the spectral curve),
$\gamma_{\alpha\beta}=\frac{4\alpha\omega_{\alpha\beta}^{3}n|\langle\alpha|\mathbf{d}|\beta\rangle|^{2}}{3c^{2}e^{2}},$ (38)
where $\alpha=1,2$ and $\beta=A,B$, are the G4V natural decay rates based on Fermi’s golden rule [47]. Here, $\alpha=1/137$ is the fine structure constant, $e$ is the elementary charge. The G4V natural decay rates relate to the decay time $T_{1}$ as $\gamma=1/T_{1}$. For the SiV (SnV), $T_{1}=1.7(4.5)$ ns [65, 47], with the Debye-Waller factor $0.8(0.6)$ [1, 47]. Additionally, we define the theoretical cooperativities [66] as a function of the coupling strength $g_{kl}$, full width at half maximum of the spectral curve $\gamma_{kl}$, and full width at half maximum of the spectral curve $2\kappa$ as
$C_{kl}=\frac{|g_{kl}|^{2}}{2\kappa\gamma_{k}}.$ (39)
In [56] the experimental cooperativity is used which is $C_{kl,\rm bh}=4C_{kl}$.
In the rotating frame after the unitary transformation of the unitary operator $U(t)=e^{-{\rm i}H_{0}t}$, where $H_{0}=\sum_{m=2}^{8}\epsilon_{m}|m\rangle\langle m|+\omega_{c}a_{c}^{\dagger}a_{c}$, the Hamiltonian can be expressed as
$\tilde{H}(t)=$ $\sum_{k\in G,l\in E}\Big(e^{{\rm i}(\epsilon_{k}-\epsilon_{l}-\omega_{c})t}g_{kl}\sigma_{kl}a+e^{{\rm i}(\epsilon_{k}-\epsilon_{l}+\omega_{c})t}g_{kl}\sigma_{kl}a^{\dagger}+e^{-{\rm i}(\epsilon_{k}-\epsilon_{l}+\omega_{c})t}g_{kl}^{*}\sigma_{lk}a+e^{-{\rm i}(\epsilon_{k}-\epsilon_{l}-\omega_{c})t}g_{kl}^{*}\sigma_{lk}a^{\dagger}\Big)\;.$ (40)
Neglecting the terms which fulfill $\epsilon_{k}-\epsilon_{l}\pm\omega_{c}\gg{\rm max}|\bm{g}|$ (rotating-wave approximation) and defining $\delta=\epsilon_{A}-\omega_{c}$ and $\omega_{s}=\epsilon_{2}$, the Hamilitonian becomes
$H_{\rm RWA}(t)=$ $e^{-{\rm i}\delta t}g_{1A}\sigma_{1A}a^{\dagger}+e^{{\rm i}(\omega_{s}-\delta)t}g_{2A}\sigma_{2A}a^{\dagger}+e^{{\rm i}(\epsilon_{A}-\epsilon_{B}-\delta)t}g_{1B}\sigma_{1B}a^{\dagger}+e^{{\rm i}(\omega_{s}-\epsilon_{B}+\epsilon_{A}-\delta)t}g_{2B}\sigma_{2B}a^{\dagger}+{\rm H.c.}$
In the frame after another unitary transformation of the unitary operator
$U(t)=e^{-{\rm i}(\delta_{A}|A\rangle\langle A|+\delta_{B}|B\rangle\langle B|)t}$ (41)
with $\delta_{A}=-\delta$ and $\delta_{B}=\omega_{s}-\epsilon_{B}+\epsilon_{A}-\delta=-\omega_{2B}+\omega_{c}$, the Hamiltonian becomes
$H_{r}(t)$ $=U^{\dagger}(t)H_{\rm RWA}(t)U(t)-{\rm i}\dot{U}(t)U(t)^{\dagger}$ (42)
$=H^{\prime}_{0}+g_{1A}\sigma_{1A}a^{\dagger}+e^{{\rm i}\omega_{s}t}g_{2A}\sigma_{2A}a^{\dagger}+e^{-{\rm i}\omega_{s}t}g_{1B}\sigma_{1B}a^{\dagger}+g_{2B}\sigma_{2B}a^{\dagger}+{\rm H.c.}\;,$ (43)
where $H^{\prime}_{0}=-\delta_{A}|A\rangle\langle A|-\delta_{B}|B\rangle\langle B|$. Now consider an input mode $a_{\rm in}$. The driving Hamiltonian is
$H_{\rm drive}(t)={\rm i}\sqrt{2k_{l}}(a_{\rm in}(t)a^{\dagger}-a_{\rm in}^{*}(t)a)$ (44)
For an open quantum system with Hamiltonian $H_{r}$ and Lindblad operators $\{L_{i}\}$, and in the Heisenberg picture, the dynamics of an operator $X$ is governed by the Heisenberg-Langevin equation
${\dot{X}}={i}[H_{r}+H_{\rm drive}(t),X]+\sum_{m=1}^{5}L^{\dagger}_{m}XL_{m}-{\frac{1}{2}}\left\{L_{m}^{\dagger}L_{m},X\right\}$ (45)
for the operators $X=a,\sigma_{1A},\sigma_{2A},\sigma_{1B},\sigma_{2B},\sigma_{11},\sigma_{22},\sigma_{AA},\sigma_{BB}$. We end up evaluating the commutators of the equation system
$\dot{a}$ $=-{\rm i}[a,H_{{r}}]-\kappa a+\sqrt{2\kappa_{l}}a_{{\rm in}}(t),$ (46)
$\dot{\sigma}_{1A}$ $=-{\rm i}[\sigma_{1A},H_{{r}}]-\frac{1}{2}(\gamma_{1A}+\gamma_{2A})\sigma_{1A},$ (47)
$\dot{\sigma}_{2A}$ $=-{\rm i}[\sigma_{2A},H_{{r}}]-\frac{1}{2}(\gamma_{1A}+\gamma_{2A})\sigma_{2A},$ (48)
$\dot{\sigma}_{1B}$ $=-{\rm i}[\sigma_{1B},H_{{r}}]-\frac{1}{2}(\gamma_{1B}+\gamma_{2B})\sigma_{1B},$ (49)
$\dot{\sigma}_{2B}$ $=-{\rm i}[\sigma_{2B},H_{{r}}]-\frac{1}{2}(\gamma_{1B}+\gamma_{2B})\sigma_{2B},$ (50)
$\dot{\sigma}_{11}$ $=-{\rm i}[\sigma_{11},H_{{r}}]+\gamma_{1A}\sigma_{AA}+\gamma_{1B}\sigma_{BB},$ (51)
$\dot{\sigma}_{22}$ $=-{\rm i}[\sigma_{22},H_{{r}}]+\gamma_{2A}\sigma_{AA}+\gamma_{2B}\sigma_{BB},$ (52)
$\dot{\sigma}_{AA}$ $=-{\rm i}[\sigma_{AA},H_{{r}}]-(\gamma_{1A}+\gamma_{2A})\sigma_{AA},$ (53)
$\dot{\sigma}_{11}+$ $\dot{\sigma}_{22}+\dot{\sigma}_{AA}+\dot{\sigma}_{BB}=0.$ (54)
For compact notation, we remove the symbol $\langle\cdot\rangle$. Due to the assumption of a single-sided cavity, it holds $\kappa_{l}=\kappa$. The Heisenberg-Langevin equations are now a nonlinear and time-varying dynamical system. They read
$\dot{a}=-{\rm i}(g_{1A}\sigma_{1A}+e^{{\rm i}\omega_{s}t}g_{2A}\sigma_{2A}+e^{-{\rm i}\omega_{s}t}g_{1B}\sigma_{1B}+g_{2B}\sigma_{2B})-\kappa a+\sqrt{2\kappa}a_{\rm in},$ (55)
$\dot{\sigma}_{1A}=-{\rm i}(\delta\sigma_{1A}+e^{-{\rm i}\omega_{s}t}g_{2A}^{*}\sigma_{1A}\sigma_{2A}^{\dagger}a-e^{{\rm i}\omega_{s}t}g_{1B}^{*}\sigma_{1B}^{\dagger}\sigma_{1A}a+g_{1A}^{*}a(\sigma_{11}-\sigma_{AA}))-\frac{1}{2}(\gamma_{1A}+\gamma_{2A})\sigma_{1A},$ (56)
$\dot{\sigma}_{2A}=-{\rm i}(\delta\sigma_{2A}+g_{1A}^{*}\sigma_{2A}\sigma_{1A}^{\dagger}a-g_{2B}^{*}\sigma_{2B}^{\dagger}\sigma_{2A}a+e^{-{\rm i}\omega_{s}t}g_{2A}^{*}a(\sigma_{22}-\sigma_{AA}))-\frac{1}{2}(\gamma_{1A}+\gamma_{2A})\sigma_{2A},$ (57)
$\dot{\sigma}_{1B}=-{\rm i}(-(\omega_{s}-\epsilon_{B}+\epsilon_{A}-\delta)\sigma_{1B}+g_{1A}^{*}\sigma_{1A}^{\dagger}\sigma_{1B}a+g_{2B}^{*}\sigma_{1B}\sigma_{2B}^{\dagger}a+e^{{\rm i}\omega_{s}t}g_{1B}^{*}a(\sigma_{11}-\sigma_{BB}))-\frac{1}{2}(\gamma_{1B}+\gamma_{2B})\sigma_{1B},$ (58)
$\dot{\sigma}_{2B}=-{\rm i}(-(\omega_{s}-\epsilon_{B}+\epsilon_{A}-\delta)\sigma_{2B}-e^{-{\rm i}\omega_{s}t}g_{2A}^{*}\sigma_{2A}^{\dagger}\sigma_{2B}a+e^{{\rm i}\omega_{s}t}\sigma_{2B}\sigma_{1B}^{\dagger}g_{1B}^{*}a+g_{2B}^{*}a(\sigma_{22}-\sigma_{BB}))-\frac{1}{2}(\gamma_{1B}+\gamma_{2B})\sigma_{2B},$ (59)
$\dot{\sigma}_{11}=-{\rm i}(g_{1A}\sigma_{1A}a^{\dagger}-g_{1A}^{*}\sigma_{1A}^{\dagger}a+e^{-{\rm i}\omega_{s}t}g_{1B}\sigma_{1B}a^{\dagger}-e^{{\rm i}\omega_{s}t}g_{1B}^{*}\sigma_{1B}^{\dagger}a)+\gamma_{1A}\sigma_{AA}+\gamma_{1B}\sigma_{BB},$ (60)
$\dot{\sigma}_{22}=-{\rm i}(e^{{\rm i}\omega_{s}t}g_{2A}\sigma_{2A}a^{\dagger}-e^{-{\rm i}\omega_{s}t}g_{2A}^{*}\sigma_{2A}^{\dagger}a+g_{2B}\sigma_{2B}a^{\dagger}-g_{2B}^{*}\sigma_{2B}^{\dagger}a)+\gamma_{2A}\sigma_{AA}+\gamma_{2B}\sigma_{BB},$ (61)
$\dot{\sigma}_{AA}=-{\rm i}(-g_{1A}\sigma_{1A}a^{\dagger}+g_{1A}^{*}\sigma_{1A}^{\dagger}a-e^{{\rm i}\omega_{s}t}g_{2A}\sigma_{2A}a^{\dagger}+e^{-{\rm i}\omega_{s}t}g_{2A}^{*}\sigma_{2A}^{\dagger}a)-(\gamma_{1A}+\gamma_{2A})\sigma_{AA},$ (62)
$\dot{\sigma}_{11}+\dot{\sigma}_{22}+\dot{\sigma}_{AA}+\dot{\sigma}_{BB}=0.$ (63)
Additionally, it holds the input-output relation for the cavity
$a_{{\rm out}}+a_{{\rm in}}=\sqrt{2\kappa}a\;.$ (64)
It is important to mention that $\kappa$ is a half-width at half maximum (HWHM). If it were a full-width at half maximum, the factor two would be removed. The boundary condition for the operators depends on the initial states. For example, for the system initialized in a photonic vacuum and spin state $|1\rangle$, we can evaluate the operator $\sigma_{11}(t=0)=1$ and $\sigma_{22}(t=0)=0$. Additionally, when calculating the outcoming mode for a broadband incoming photon, we assume an incoming mode $a_{\rm in}(t)=e_{0}e^{({\rm i}\omega_{0}-\gamma/2)t}$ with amplitude $e_{0}$, central frequency $\omega_{0}$ and bandwidth $\gamma$.
If the rotating wave approximation (RWA) allows for neglecting cross couplings, the spin-dependent reflection is modeled using the reflection coefficients.
$R_{1}(\omega)$ $=-1+\frac{2\kappa({\rm i}(\omega-\omega_{1A})+\gamma_{\rm avg,A})}{({\rm i}(\omega-\omega_{c})+\kappa)({\rm i}(\omega-\omega_{1A})+\gamma_{\rm avg,A})+|g_{1A}|^{2}},$ (65)
$R_{2}(\omega)$ $=-1+\frac{2\kappa({\rm i}(\omega-\omega_{2B})+\gamma_{\rm avg,B})}{({\rm i}(\omega-\omega_{c})+\kappa)({\rm i}(\omega-\omega_{2B})+\gamma_{\rm avg,B})+|g_{2B}|^{2}},$ (66)
with $\gamma_{\rm avg,A}=\frac{1}{2}(\gamma_{1A}+\gamma_{2A})$ and $\gamma_{\rm avg,B}=\frac{1}{2}(\gamma_{1B}+\gamma_{2B})$. Since $\gamma_{1A}\gg\gamma_{2A}$ and $\gamma_{2B}\gg\gamma_{1B}$ it is sufficient to set $\gamma_{\rm avg,A}=\frac{1}{2}\gamma_{1A}$ and $\gamma_{\rm avg,B}=\frac{1}{2}\gamma_{2B}$. Previous studies derived the reflection spectrum of a three-level atom [55] and G4V [49]. It is important to point out that $\kappa$ is a half-width at half maximum. In [55], the atomic decay rate $\gamma:=\gamma_{1A}/2$ is used, meaning that $\gamma$ is a half-width at half maximum as well. Note that $\kappa$ and $\gamma$ can be defined either as the half-width at half maximum (HWHM) or the full-width at half maximum (FWHM); care should be taken to ensure consistency.
In [49], the assumed magnetic field is assumed to be close to the defect’s symmetry axis such that the coupling strengths $g_{2A},g_{1B}$ are small enough to neglect them. In our modeling, we include the cross-talk effects and solve the system of equations above with numerical methods shown in [58] for solving ordinary differential equations.
APPENDIX B OPTIMIZING SPIN-PHOTON CPHASE GATES
For the optimization of the controlled phase gate we neglect cross-talk. To derive the objective function we firstly elaborate on the reflection scheme for broadband incoming photons. We model broadband incoming photons in time-domain by
$a(t)=\epsilon_{0}e^{({\rm i}\omega_{0}-\gamma/2)t}.$ (67)
The amplitude $\epsilon_{0}$ is so small such that no driving between the ground and excited state is steered. The spectrum is given by
$\tilde{S}(\omega-\omega_{0})=\frac{\epsilon_{0}}{{\rm i}(\omega-\omega_{0})+\gamma/2}.$ (68)
$\begin{split}|\psi_{{\rm ph}}\rangle|{1}\rangle=&\int_{\mathbb{R}}\tilde{S}(\omega-\omega_{0})\left(\alpha|\omega\rangle_{e}|{1}\rangle+\beta|\omega\rangle_{l}|{1}\rangle\right){\rm d}\omega\\ \overset{{\rm early\,reflection}}{\rightarrow}&\int_{\mathbb{R}}\tilde{S}(\omega-\omega_{0})\left(R_{1}(\omega)\alpha|\omega\rangle_{e}|{1}\rangle+\beta|\omega\rangle_{l}|{1}\rangle\right){\rm d}\omega\\ \overset{\pi/2\,{\rm rotation}}{\rightarrow}&\int_{\mathbb{R}}\tilde{S}(\omega-\omega_{0})\left(R_{1}(\omega)\frac{\alpha}{\sqrt{2}}|\omega\rangle_{e}(|{1}\rangle+|{2}\rangle)+\frac{\beta}{\sqrt{2}}|\omega\rangle_{l}(|{1}\rangle+|{2}\rangle)\right){\rm d}\omega\\ \overset{{\rm late\,reflection}}{\rightarrow}&\int_{\mathbb{R}}\tilde{S}(\omega-\omega_{0})\Bigg(R_{1}(\omega)\frac{\alpha}{\sqrt{2}}|\omega\rangle_{e}(|{1}\rangle+|{2}\rangle)\\ +&\frac{\beta}{\sqrt{2}}|\omega\rangle_{l}(R_{1}(\omega)|{1}\rangle+R_{2}(\omega)|{2}\rangle)\Bigg){\rm d}\omega\end{split}$ (69)
Assuming that every photon gets detected with unity probability the $X$-measurement has the form [67]
$\rho_{+}=\int{}_{\omega}\langle+|\psi\rangle\langle\psi|+\rangle_{\omega}\,{\rm d}\omega,$ (70)
$\rho_{-}=\int{}_{\omega}\langle-|\psi\rangle\langle\psi|-\rangle_{\omega}\,{\rm d}\omega.$ (71)
The respective entries are
$\langle 1|\rho_{\pm}|1\rangle$ $=\frac{1}{4}|\alpha\pm\beta\rvert^{2}I_{1},$ (72)
$\langle 1|\rho_{\pm}|2\rangle$ $=\frac{1}{4}\alpha^{*}(\alpha\pm\beta)I_{1}+\frac{1}{4}\beta^{*}(\alpha\pm\beta)I_{2},$ (73)
$\langle 1|\rho_{\pm}|2\rangle$ $=\langle 2|\rho_{\pm}|1\rangle^{*},$ (74)
$\langle 2|\rho_{\pm}|2\rangle$ $=\frac{1}{4}|\alpha|^{2}I_{1}\pm\frac{1}{4}\alpha^{*}\beta I_{2}^{*}\pm\frac{1}{4}\alpha\beta^{*}I_{2}+\frac{1}{4}|\beta|^{2}I_{3}$ (75)
with the integrals
$I_{1}$ $=\int_{\mathbb{R}}S(\omega-\omega_{0})|R_{1}(\omega)|^{2}\,{\rm d}\omega,$ (76)
$I_{2}$ $=\int_{\mathbb{R}}S(\omega-\omega_{0})R_{1}(\omega)R_{2}^{*}(\omega)\,{\rm d}\omega,$ (77)
$I_{3}$ $=\int_{\mathbb{R}}S(\omega-\omega_{0})|R_{2}(\omega)|^{2}\,{\rm d}\omega.$ (78)
and $S(\omega)=\gamma/(2\pi e_{0}^{2})|\tilde{S}(\omega)|^{2}$. The total spin state reads
$\rho=R_{y}(\pi/2)\rho_{+}R_{y}^{\dagger}(\pi/2)+\sigma_{z}R_{y}(\pi/2)\rho_{-}R_{y}^{\dagger}(\pi/2)\sigma_{z}$ (79)
with the $\pi/2$ rotation $R_{y}(\pi/2)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$ and the Pauli matrix $\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. We apply the $\pi/2$ rotation for getting the state
$\rho_{\rm tgt}=|\psi\rangle\langle\psi|$ (80)
with
$|\psi\rangle=\alpha|2\rangle+\beta|1\rangle$ (81)
as the target state. We define $\eta_{\rm sp}={\rm tr}(\tilde{\rho}_{\rm sp})$ as the success probability of spin-photon entanglement and $\rho_{\rm sp}=\tilde{\rho}_{\rm sp}/\eta_{\rm sp}$ as the state. We optimize for cavity parameters $(\kappa,\omega_{c})$ and emission central frequency $\omega_{0}$ such that the fidelity
$F_{\rm sp}=\langle\rm Bell|\rho_{\rm sp}|\rm Bell\rangle$ (82)
with $|\rm Bell\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ is maximized for $\alpha=\beta=1/\sqrt{2}$.
APPENDIX C READ-IN PROCESS
For the read-in process we model a photon source emitting broadband photons of bandwidth $\gamma$ and fidelity $F_{\rm ph}$ and assume an imperfect $\pi/2$ spin rotation produced by either microwave or optical spin control. Starting point is the photonic qubit
$|\psi_{0}\rangle=\frac{1}{\sqrt{2}}(|E\rangle+|L\rangle).$ (83)
To model imperfect photon pair generation we apply a depolarizing channel [68], i.e.
$\rho=(1-\epsilon)\rho_{0}+\epsilon I/2$ (84)
with $\epsilon=2(1-F_{\rm ph})$, the fidelity $F_{\rm ph}$ of the photon pair generated by the quantum dot and the pure state $\rho_{0}=|\psi_{0}\rangle\langle\psi_{0}|$. For simplicity we employ the notation
$\rho=\sum\rho_{IJKM}|IJ\rangle\langle KM|$ (85)
with $I,K=E,L$. We assume an imperfect spin $\pi/2$ rotation modeled by a map $\Lambda$, i.e.
$\mathcal{D}_{\pi/2}(|1\rangle\langle 1|)=\sum\Lambda^{(mk)}|m\rangle\langle k|.$ (86)
The reflection scheme entails the early reflection, a spin $\pi/2$ rotation and a late reflection and reads
$\sum\rho_{IK}|I\rangle\langle K|1\rangle\langle 1|=\sum\rho_{IK}a^{I}\bar{a}^{K}|1\rangle\langle 1|$ (87)
$\xrightarrow[]{\text{early reflection}}$ $\sum\rho_{IK}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}a^{L})$ (88)
$(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{a}^{L})|1\rangle\langle 1|$ (89)
$\xrightarrow[]{\pi/2\,\,\text{rotation}}$ $\sum\rho_{IK}\sum\Lambda^{(mk)}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}a^{L})$ (90)
$(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{a}^{L})|m\rangle\langle k|$ (91)
$\xrightarrow[]{\text{late reflection}}$ $\sum\rho_{IK}\sum\Lambda^{(mk)}(\delta_{IE}\mathcal{D}_{1}(a^{E})+\delta_{IL}\mathcal{D}_{m}(a^{L}))$ (92)
$(\delta_{KE}\bar{\mathcal{D}}_{1}(a^{E})+\delta_{KL}\bar{\mathcal{D}}_{k}(\bar{a}^{L}))|m\rangle\langle k|$ (93)
where $\mathcal{D}_{k}(a)$ denotes the outcoming mode $a_{\rm out}$ from the time evolution according to the Heisenberg-Langevin equations shown in Eqs. (55)-(64) initializing $\sigma_{kk}(0)=1$ for an input mode $a_{\rm in}$. After the $X$-measurement the state is
$\rho_{\pm}=\frac{1}{2}\sum_{IK}\sum_{m,k}\Lambda^{(mk)}\left(\delta_{IE}\delta_{KE}I_{1}\pm\delta_{IE}\delta_{KL}I_{k}\pm\delta_{IL}\delta_{KE}I_{2m-1}+\delta_{IL}\delta_{KL}I_{2(m-1)+k}\right)$ (94)
with $\mathbf{I}=(I_{1},I_{2},I_{2}^{*},I_{3})$ and
$I_{1}=\mathcal{N}\int_{0}^{T}|\mathcal{D}_{1}(a_{\rm in}(t))|^{2}\,{\rm d}t,$ (95)
$I_{2}=\mathcal{N}\int_{0}^{T}\mathcal{D}_{1}(a_{\rm in}(t))\mathcal{D}_{2}^{*}(a_{\rm in}(t))\,{\rm d}t,$ (96)
$I_{3}=\mathcal{N}\int_{0}^{T}|\mathcal{D}_{2}(a_{\rm in}(t))|^{2}\,{\rm d}t$ (97)
with $\mathcal{N}=\gamma/e_{0}^{2}$ and $T$ sufficiently large.
APPENDIX D READ-OUT PROCESS
For the read-out process we assume a perfect CPHASE gate between the spin and photon and no depolarization during photon generation for simplicity. The process requires an extra single photon source emitting single photons of the form $|\psi_{\rm ph}\rangle=1/\sqrt{2}(|e\rangle+|l\rangle)$ with the corresponding density matrix $\rho_{\rm ph}=|\psi_{\rm ph}\rangle\langle\psi_{\rm ph}|$. For the computation of Kraus-operators for the read-out process we apply the reflection scheme on the full basis of $2\times 2$ matrices for the spin. The initial state is $\rho^{ij}=\rho_{\rm ph}\otimes|i\rangle\langle j|$. The controlled phase gates applied here are $U_{e}={\rm diag}(-1,1,1,1)$ and $U_{l}={\rm diag}(1,1,-1,1)$. We assume an imperfect $\pi/2$ rotation for the spin qubit driven by either optical or microwave control modeled by the Kraus-operators $\{K_{m}\}$. After the late reflection the state reads
$\rho_{\rm lr}^{ij}=U_{l}\left(\sum_{m=1}^{4}K_{m}\left(U_{e}\rho^{ij}U_{e}^{\dagger}\right)K_{m}^{\dagger}\right)U_{l}^{\dagger},\quad i,j=1,2$ (98)
The measurement outcomes in the $Z$-basis are $\rho_{11}^{ij}=\langle 1|\rho^{ij}|1\rangle$ and $\rho_{22}^{ij}=\langle 2|\rho^{ij}|2\rangle$.
Evaluating the photonic qubit for $i,j=1,2$ results in the Kraus-operators for the read-out process.
APPENDIX E KRAUS-OPERATORS
E.1 Optical Control
$K_{1}^{\rm in}$ $=\begin{pmatrix}0.527&0.527\\ 0.385+0.015\,{\rm i}&-0.379-0.013\,{\rm i}\end{pmatrix}$ (99)
$K_{2}^{\rm in}$ $=\begin{pmatrix}0.152-0.005\,{\rm i}&0.152-0.005\,{\rm i}\\ -0.204-0.001\,{\rm i}&0.215\end{pmatrix}$ (100)
$K_{3}^{\rm in}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (101)
$K_{4}^{\rm in}$ $=\begin{pmatrix}-0.001&-0.001\\ 0.08&0.078\end{pmatrix}$ (102)
$K_{1}^{\rm out}$ $=\begin{pmatrix}0.51&0.417-0.022\,{\rm i}\\ 0.51&-0.417+0.022\,{\rm i}\end{pmatrix}$ (103)
$K_{2}^{\rm out}$ $=\begin{pmatrix}-0.205-0.011\,{\rm i}&0.251\\ -0.205-0.011\,{\rm i}&-0.251\end{pmatrix}$ (104)
$K_{3}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (105)
$K_{4}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (106)
E.2 Microwave Control
$K_{1}^{\rm in}$ $=\begin{pmatrix}0.446&0.446\\ 0.442+0.024\,{\rm i}&-0.301+0.244\,{\rm i}\end{pmatrix}$ (107)
$K_{2}^{\rm in}$ $=\begin{pmatrix}0.045+0.037\,{\rm i}&0.045+0.037\,{\rm i}\\ 0.050+0.044\,{\rm i}&0.210\end{pmatrix}$ (108)
$K_{3}^{\rm in}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (109)
$K_{4}^{\rm in}$ $=\begin{pmatrix}-0.027+0.001\,{\rm i}&-0.027+0.001\,{\rm i}\\ 0.053+0\,{\rm i}&-0.002+0.001\,{\rm i}\end{pmatrix}$ (110)
$K_{1}^{\rm out}$ $=\begin{pmatrix}-0.499-0.024\,{\rm i}&-0.500\\ -0.499-0.024\,{\rm i}&0.500\end{pmatrix}$ (111)
$K_{2}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (112)
$K_{3}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (113)
$K_{4}^{\rm out}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$ (114)